# Hyper-STTN: Social Group-aware Spatial-Temporal Transformer Network for Human Trajectory Prediction with Hypergraph Reasoning

### v2

@@ IV-D Discussion
Trajectory results for two specific sparse and dense cases of Hyper-STTN are visualized alongside three ablation models in the same configuration, as shown in Fig. 5. The observation input of each network is represented by blue points, while the groundturth for pedestrians is colored in green, distinguishing it from the forecasting outputs of the above predictors. It is evident that Hyper-STTN consistently generates highly accurate locations for several sequential timesteps and infers more reasonable motion orientations. As shown in Fig. 6, the developments in the spatial transformer allows Hyper-STTN to understand relative spatial interactions among each agent pair. On the other hand, the encoded embeddings from the temporal transformer lead to an understanding of individual motion attributes. Additionally, in dense scenarios with an increasing number of pedestrian groups, the diverse intergroup and intragroup crowd correlations have a more significant impact than in sparse situations. Hence, the inference of group-wise interactions results in more accurate long-term forecasting, as observed in the trajectories of Hyper-STTN and HGNN, along with their corresponding ADE/FDE results.
@@ IV-D1 Effect of Hypergraph Neural Network
We address three components of Hyper-STTN, which include the hypergraph neural network, spatial-temporal transformer, and multi-modal transformer. As shown in Table I, HGNN performs approximately as well as GroupNet on many datasets. In general, hypergraph construction can effectively enhance network performance as the number of pedestrians increases. For example, in Fig. 5d, the ablation model STTN exhibits serious forecasting errors of many agents, both in terms of location and orientation, when compared to HGNN. And the crowd hypergraph group constructions are shown in in Fig. 6 which is attributed to STTN’s lack of consideration for group-wise features.
@@ IV-D2 Effect of Spatial-Temporal Transformer
Compared to previous transformer-based approaches such as STAR, STTN improves the average results from 0.26/0.53 to 0.25/0.51. This improvement can be attributed to the deployment of a more effective transformer-based multi-modal fusion network and mask attention mechanism. Additionally, STTN also achieves more efficient trajectories than HGNN in scenarios with fewer pedestrians, as demonstrated in Fig. 5b.
IV-D3 Effect of Multi-Modal Transformer
Due to the heterogeneity of group-wise and pair-wise embeddings, Hyper-STTN leverages a multi-modal transformer to merge them, capturing cross-modal attention to highlight meaningful correlations among modalities. In contrast, the limited fusion mechanism in the ablation model Hyper-STTN+MLP underperforms Hyper-STTN in every dataset, as shown in Table I and Table II. This is because the direct concatenation operation in the MLP overlooks potential relationships among cross-modal features.
@@ IV-D3 Effect of Multi-Modal Transformer
Figure 6: The illustration of group-wise and pair-wise interactions in Hyper-STTN: The pair-wise attention maps are illustrated on subfigures (a) and (b), the red and black circles represent the attention score of current agent and neighboring agents respectively. Additionally, crowd group-wise interactions are described by hypergraphs with different scales in subfigures (c) and (d).
@@ V Conclusion
In this work, we introduced Hyper-STTN, a hypergraph-based hybrid spatial-temporal transformer for trajectory prediction tasks. Hyper-STTN captures both group-wise and pair-wise interactions in crowd dynamics using hypergraphs and transformer blocks. It has outperformed the state-of-the-art prediction algorithm on the ETH-UCY, NBA, and SDD dataset. Future work could explore further enhancements to Hyper-STTN, including refining its ability to handle complex scenarios and scaling it for real-time applications.

### v3

@@ IV-D Discussion
Fig. 5 illustrates the trajectory predictions of Hyper-STTN under both sparse and dense scenarios, alongside two ablation models and an additional variant that denotes the multimodal transformer is replaced by an MLP under otherwise identical settings. In all visualizations, blue dots indicate the observation inputs, while green dots denote the pedestrian ground truth trajectories, allowing clear differentiation from the predicted trajectories. As shown, Hyper-STTN consistently produces highly accurate multi step location forecasts and infers more realistic motion orientations. Fig. 6 further highlights how the spatial transformer module enables Hyper-STTN to capture relative spatial interactions among agent pairs, whereas the temporal transformer encodes individual motion attributes. Moreover, in dense scenarios with a growing number of pedestrian groups, inter-group and intra-group crowd correlations exert a stronger influence than in sparse conditions. Consequently, explicitly inferring groupwise interactions yields more accurate long-term forecasts, as evidenced by the trajectories and ADE/FDE metrics achieved by Hyper-STTN and HGNN.
@@ IV-D1 Effect of Hypergraph Neural Network
Hyper-STTN comprises two key components: a hypergraph neural network, a spatio-temporal transformer with the fusion model. As reported in Table. I, the HGNN module achieves performance comparable to GroupNet on several datasets, demonstrating that hypergraph construction effectively enhances prediction accuracy as crowd density increases. For instance, Fig. 5d shows that the ablation model STTN incurs substantial forecasting errors in both agent locations and orientations compared with HGNN. The constructed crowd hypergraphs visualized in Fig. 6 further highlight that these improvements stem from STTN’s lack of explicit modeling of groupwise features, which HGNN introduces.
@@ IV-D2 Effect of Spatial-Temporal Transformer
Compared to previous transformer-based approaches such as STAR, STTN improves the average results from 0.26/0.53 to 0.25/0.51. This improvement can be attributed to the deployment of a more effective transformer-based multi-modal fusion network and mask attention mechanism. Additionally, STTN also achieves more efficient trajectories than HGNN in scenarios with fewer pedestrians, as demonstrated in Fig. 5b. More experimental visualizations and code illustrations can be found at the project website: https://sites.google.com/view/hypersttn.
@@
Figure 6: The illustration of groupwise and pairwise interactions in Hyper-STTN: The pairwise attention maps are illustrated on subfigures (a) and (b), the red and black circles represent the attention score of current agent and neighboring agents respectively. Additionally, crowd groupwise interactions are described by hypergraphs with different scales in subfigures (c) and (d).
@@ V Conclusion
In this paper, we proposed Hyper-STTN, a hypergraph-based hybrid spatial-temporal transformer for trajectory prediction tasks. Hyper-STTN jointly captures groupwise and pairwise interactions in crowd dynamics through hypergraph and transformer networks. Extensive experiments on the ETH-UCY and NBA datasets demonstrate that Hyper-STTN outperforms existing state-of-the-art prediction algorithms. Future research will focus on enhancing its capability to handle more complex scenarios and on scaling the model for real-time deployments.